# Astrometric properties of reference frame sources as a function of redshift

### v1

@@ 3.4 Photometric properties
The main differences between different types of RFC sources, based on their spectral classification, are most clearly seen in the relationship between magnitude and redshift. As shown in Fig. 6, we examine the relationships between redshift and the three Gaia magnitudes ($G$, BPBP, and $RP$) of the sample. We find that the median $G$ magnitude of galaxies remains roughly constant at $\sim\!20$ and shows no significant dependence on redshift, while the median $G$ magnitude of QSOs is around $\sim\!19$ at the same redshift. Therefore, although the $G$ magnitudes of both galaxies and QSOs show no strong redshift dependence individually, the concentration of galaxies within 0<z<10\!<z\!<1 combined with their systematically fainter $G$ magnitudes compared to QSOs results in a noticeable negative gradient in the median $G$ magnitude of the total sample over this redshift range. In contrast, the $BP$ and $RP$ magnitudes of galaxies exhibit a strong dependence on redshift. Within 0<z<10\!<z\!<1, the median $BP$ magnitude of galaxies increases sharply from $\sim\!17$ at z=0z\!=\!0 to $\sim\!20.5$ at z=1z\!=\!1, while the median $RP$ magnitude increases from $\sim\!16$ to $\sim\!19.5$ over the same redshift range. However, QSOs exhibit only a slight increase with increasing redshift in $G$, BPBP and $RP$ magnitudes. This difference is reflected in the median $BP$ and $RP$ magnitudes of the total sample. They increase rapidly over 0<z<0.50\!<z\!<0.5, followed by a slower increase over 0.5<z<10.5\!<z\!<1. However, at z>1z\!>\!1, the sample is dominated by QSOs and the median magnitudes show only a weak increase with redshift.
@@
The relationships between $G$, BPBP, and $RP$ magnitudes and redshift also influence evolution of the corresponding colour indices with redshift. As shown in Fig. 7, we examine the relationships between redshift and three colour indices: BPRPBP-RP (left), GRPG-RP (middle), and BPGBP-G (right). These colour indices exhibit relatively tighter distributions compared to the magnitudes. We find that the median BPRPBP-RP colour of the total sample decreases sharply with increasing redshift over 0<z<10\!<z\!<1, remains nearly constant at $\sim\!0.6$ over 1<z<31\!<z\!<3 (with a slight increase to $\sim\!0.7$ near $z\!\simeq\!1.5$), and rises significantly at z>3z\!>\!3. This trend is consistent with that of the QSOs. For galaxies, the median BPRPBP-RP remains around $\sim\!1.5$ within 0<z<0.50\!<z\!<0.5, which likely contributes to the negative gradient observed in the total sample over this redshift range. The redshift dependence of the median GRPG-RP and BPGBP-G colours is mainly determined by the strong redshift dependence of the median $BP$ and $RP$ magnitudes in galaxies, whereas the median $G$ magnitude remains nearly constant over 0<z<10\!<z\!<1. As a result, the median GRPG-RP colour of galaxies decreases sharply and the median BPGBP-G colour increases sharply with redshift within this range. For z>1z\!>\!1, the median GRPG-RP colour of QSOs is concentrated around $\sim\!0.4$ while the median BPGBP-G colour is concentrated around zero, and neither of them shows significant dependence on redshift. Since galaxies dominate the sample at low redshift (0<z<0.50\!<z\!<0.5), the total sample exhibits strong gradients in the median GRPG-RP and BPGBP-G colours over this range, consistent with the trends of galaxies. Over 0.5<z<10.5\!<z\!<1, the fraction of galaxies decreases and the evolution becomes more gradual. At higher redshift (z>1z\!>\!1), the sample is dominated by QSOs and the median values remain nearly constant, at $\sim\!0.4$ and $\sim\!0.1$, respectively. It is worth noting that the median GRPG-RP colour of QSOs shows a slight enhancement near $z\!\simeq\!1.5$ and a modest increase at z>3z\!>\!3, consistent with the trend of the median BPRPBP-RP colour with redshift.
@@ 4.1 Comparison with existing redshift catalogues
In the RFC cross-match table, redshift information is available for approximately 8 900 sources, most of which were compiled from NED. These redshifts originate from a mixture of spectroscopic and photometric measurements, as well as machine-learning estimates. About 2 500 sources (28%) are based on SDSS DR13 spectroscopy (Albareti et al. 2017), while an additional 713 sources (8%) rely on photometric redshifts from SDSS DR6 (Richards et al. 2009) and the Two Micron All Sky Survey (2MASS; Skrutskie et al. 2006), which are subject to relatively large uncertainties. In this work, we substantially expand and update the spectroscopic redshift content by incorporating the latest measurements from SDSS DR17 and DR19, and in particular from DESI DR1 (released in March 2025; DESI Collaboration et al. 2025). As a result, we obtain a sample of 4 774 RFC sources with high-quality spectroscopic redshifts, accounting for more than 50% of the redshifts in the RFC cross-match table. During the compilation process, we carefully considered redshift quality indicators (e.g. ZWARN and ZWARNING) and performed cross-checks between independent measurements from SDSS and DESI. This procedure ensures a highly reliable spectroscopic redshift sample, which forms the basis of our analysis of redshift-dependent astrometric properties.
@@
The RFC cross-match table also provides source classifications primarily based on NED. We refine these classifications using spectroscopic information from SDSS and DESI, adopting a simplified scheme that separates RFC sources into galaxies and QSOs, while excluding objects with uncertain or missing classifications. This approach reduces classification complexity and facilitates a clearer comparison of redshift-dependent astrometric behaviour between the two populations. In future work, we plan to implement a more detailed spectroscopic classification, similar to that of Sexton et al. (2022), by identifying subclasses such as Seyfert I, Seyfert II, and blazars, and to include this information in an updated RFC redshift catalogue.
@@ 4.3 Difference between QSOs and galaxies
In this work, we separate RFC sources into galaxies and QSOs to investigate differences in the redshift evolution of their astrometric properties. At the same redshift, galaxies are systematically fainter than QSOs, with median $G$ magnitude of $\sim\!20$ and $\sim\!19$, respectively (Fig. 6). Since the $G$ magnitude is the dominant factor determining Gaia astrometric precision (Gaia Collaboration et al. 2023), this systematic faintness leads galaxies to exhibit larger optical astrometric uncertainties and consequently larger radio-optical positional offsets. In addition, galaxies show more negative measured parallaxes and larger proper motions than QSOs at the same redshift, reflecting increased measurement noise rather than intrinsic motions.
@@ 5 Conclusion
We have combined DESI DR1 and SDSS DR17/DR19 spectroscopy with Gaia DR3 to construct a large and homogeneous sample of RFC sources with reliable redshifts and classifications. Using this sample, we investigated the redshift dependence of astrometric parameters from VLBI observations and Gaia DR3, providing additional observational support for the consistency between the optical and radio celestial reference frames.
@@ Acknowledgements.
We sincerely thank the anonymous referee, Sébastien Lambert, Leonid Petrov, Jun Yang, Xiaopeng Cheng and Bo Zhang for the fruitful discussions and useful suggestions, which have greatly improved the work. This work was supported by the National Natural Science Foundation of China (NSFC) under grant Nos. 12573070 and 12373074. We acknowledge the use of the Radio Fundamental Catalogue (Petrov and Kovalev 2025). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations are managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research used material from the Bordeaux VLBI Image Database (BVID). This database can be reached at http://bvid.astrophy.u-bordeaux.fr/. We made much use of Astropy (http://www.astropy.org) – a community-developed core Python package for astronomy (Astropy Collaboration et al. 2018), the Python 2D plotting library Matplotlib (Hunter 2007) and TOPCAT (Taylor 2011). This work benefited from the LaTeX service provided by the e-Science Center of the Collaborative Innovation Center of Advanced Microstructures, Nanjing University (https://sci.nju.edu.cn). Researchers interested in the sample-related data may contact us via email.
@@ Appendix B Cross-checking of redshifts from different methods
As shown in Fig. 9,we perform a cross-checking of redshifts between different methods. DESI spectroscopic redshifts are compared with the literature-based redshifts (panel a), and SDSS spectroscopic redshifts from DR17 and DR19 (panel c). Similarly, SDSS spectroscopic redshifts are cross-checked against the literature-based redshifts (panel b). We find that the majority of sources show good agreement in redshifts across different methods ($|\Delta z|\!<\!0.1$). However, we find that DESI spectroscopic redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 show significant deviations from redshifts measured by other methods (panels a and c). In this work, for sources with both SDSS and DESI spectroscopic redshifts, we retain only those with consistent redshifts satisfying $|\Delta z|\!<\!0.1$ (panel d).
@@ Appendix C Astrometric properties as a function of $G$ magnitude
In this work, we find that the absolute radio-optical offsets (Figs. 10), the Gaia optical positional uncertainties (Figs. 11), parallax uncertainties (12), proper motion and the proper motion uncertainties (13) in our sample all increase steeply toward fainter $G$ magnitudes and can be well described by an exponential dependence on $G$. In contrast, as shown in Fig. 11, the median radio positional uncertainties increase only mildly and approximately linearly with $G$ magnitude, from about 0.5 mas at $G\!\sim\!18$ to roughly 1.5 mas at $G\!\sim\!21$, which is significantly smaller than the corresponding optical positional uncertainties at faint $G$ magnitudes. However, at a given $G$ magnitude, both the absolute and relative radio–optical offsets are significantly larger for galaxies than for QSOs. The Gaia optical positional uncertainties are also significantly larger for galaxies than for QSOs, while the radio positional uncertainties are comparable between galaxies and QSOs. Taken together, these results indicate that the observed radio–optical offsets of RFC sources are most likely dominated by uncertainties in the Gaia optical positions rather than in the VLBI radio measurements.

### v2

@@ 3.4 Photometric properties
The main differences between different types of RFC sources, based on their spectral classification, are most clearly seen in the relationship between magnitude and redshift. As shown in Fig. 6, we examined the relationships between redshift and the three Gaia magnitudes ($G$, BPBP, and $RP$) of the sample. We find that the median $G$ magnitude of galaxies remains roughly constant at $\sim\!20$ and shows no significant dependence on redshift, while the median $G$ magnitude of QSOs is around $\sim\!19$ at the same redshift. Therefore, although the $G$ magnitudes of both galaxies and QSOs show no strong redshift dependence individually, the concentration of galaxies within 0<z<10\!<z\!<1 combined with their systematically fainter $G$ magnitudes compared to QSOs results in a noticeable negative gradient in the median $G$ magnitude of the total sample over this redshift range. In contrast, the $BP$ and $RP$ magnitudes of galaxies exhibit a strong dependence on redshift. Within 0<z<10\!<z\!<1, the median $BP$ magnitude of galaxies increases sharply from $\sim\!17$ at z=0z\!=\!0 to $\sim\!20.5$ at z=1z\!=\!1, while the median $RP$ magnitude increases from $\sim\!16$ to $\sim\!19.5$ over the same redshift range. However, QSOs exhibit only a slight increase with increasing redshift in $G$, BPBP and $RP$ magnitudes. This difference is reflected in the median $BP$ and $RP$ magnitudes of the total sample. They increase rapidly over 0<z<0.50\!<z\!<0.5, followed by a slower increase over 0.5<z<10.5\!<z\!<1. However, at z>1z\!>\!1, the sample is dominated by QSOs and the median magnitudes show only a weak increase with redshift.
@@
The relationships between $G$, BPBP, and $RP$ magnitudes and redshift also influence the evolution of the corresponding colour indices with redshift. As shown in Fig. 7, we examined the relationships between redshift and three colour indices: BPRPBP-RP (left), GRPG-RP (middle), and BPGBP-G (right). These colour indices exhibit relatively tighter distributions compared to the magnitudes. We find that the median BPRPBP-RP colour of the total sample decreases sharply with increasing redshift over 0<z<10\!<z\!<1, remains nearly constant at $\sim\!0.6$ over 1<z<31\!<z\!<3 (with a slight increase to $\sim\!0.7$ near $z\!\simeq\!1.5$), and rises significantly at z>3z\!>\!3. This trend is consistent with that of the QSOs. For galaxies, the median BPRPBP-RP remains around $\sim\!1.5$ within 0<z<0.50\!<z\!<0.5, which likely contributes to the negative gradient observed in the total sample over this redshift range. The redshift dependence of the median GRPG-RP and BPGBP-G colours is mainly determined by the strong redshift dependence of the median $BP$ and $RP$ magnitudes in galaxies, whereas the median $G$ magnitude remains nearly constant over 0<z<10\!<z\!<1. As a result, the median GRPG-RP colour of galaxies decreases sharply and the median BPGBP-G colour increases sharply with redshift within this range. For z>1z\!>\!1, the median GRPG-RP colour of QSOs is concentrated around $\sim\!0.4$ while the median BPGBP-G colour is concentrated around zero, and neither of them shows significant dependence on redshift. Since galaxies dominate the sample at low redshift (0<z<0.50\!<z\!<0.5), the total sample exhibits strong gradients in the median GRPG-RP and BPGBP-G colours over this range, consistent with the trends of galaxies. Over 0.5<z<10.5\!<z\!<1, the fraction of galaxies decreases and the evolution becomes more gradual. At higher redshift (z>1z\!>\!1), the sample is dominated by QSOs and the median values remain nearly constant, at $\sim\!0.4$ and $\sim\!0.1$, respectively. It is worth noting that the median GRPG-RP colour of QSOs shows a slight enhancement near $z\!\simeq\!1.5$ and a modest increase at z>3z\!>\!3, consistent with the trend of the median BPRPBP-RP colour with redshift.
@@ 4.1 Comparison with existing redshift catalogues
In the RFC cross-match table, redshift information is available for approximately 8 900 sources, most of which were compiled from NED. These redshifts originate from a mixture of spectroscopic and photometric measurements, as well as machine-learning estimates. About 2 500 sources (28%) are based on SDSS DR13 spectroscopy (Albareti et al. 2017), while an additional 713 sources (8%) rely on photometric redshifts from SDSS DR6 (Richards et al. 2009) and the Two Micron All Sky Survey (2MASS; Skrutskie et al. 2006), which are subject to relatively large uncertainties. In this work, we substantially expanded and updated the spectroscopic redshift content by incorporating the latest measurements from SDSS DR17 and DR19, and in particular from DESI DR1 (released in March 2025; DESI Collaboration et al. 2025). As a result, we obtain a sample of 4 774 RFC sources with reliable spectroscopic redshifts, accounting for more than 50% of the redshifts in the RFC cross-match table. During the compilation process, we carefully considered redshift quality indicators (e.g. ZWARN and ZWARNING) and performed cross-checks between independent measurements from SDSS and DESI. This procedure ensures a reliable spectroscopic redshift sample, forming the basis of our analysis of redshift-dependent astrometric properties.
@@
The RFC cross-match table also provides source classifications primarily based on NED. We refined these classifications using spectroscopic information from SDSS and DESI, adopting a simplified scheme that separates RFC sources into galaxies and QSOs, while excluding objects with uncertain or missing classifications. This approach reduces classification complexity and facilitates a clearer comparison of redshift-dependent astrometric behaviour between the two populations. In future work, we plan to implement a more detailed spectroscopic classification, similar to that of Sexton et al. (2022), by identifying subclasses such as Seyfert I, Seyfert II, and blazars, and to include this information in an updated RFC redshift catalogue.
@@ 4.3 Difference between QSOs and galaxies
In this work, we separate RFC sources into galaxies and QSOs to investigate differences in the redshift evolution of their astrometric properties. At the same redshift, galaxies are systematically fainter than QSOs, with median $G$ magnitudes of $\sim\!20$ and $\sim\!19$, respectively (Fig. 6). Since the $G$ magnitude is the dominant factor determining Gaia astrometric precision (Gaia Collaboration et al. 2023), this systematic faintness leads galaxies to exhibit larger optical astrometric uncertainties and consequently larger radio-optical positional offsets. In addition, galaxies show more negative measured parallaxes and larger proper motions than QSOs at the same redshift, reflecting increased measurement noise rather than intrinsic motions.
@@ 5 Conclusion
We have combined DESI DR1 and SDSS DR17/DR19 spectroscopy with Gaia DR3 to construct a large and homogeneous sample of RFC sources with reliable redshifts and classifications. Using this sample, we investigate the redshift dependence of astrometric parameters from VLBI observations and Gaia DR3, providing additional observational support for the consistency between the optical and radio celestial reference frames.
@@ Acknowledgements.
We sincerely thank the anonymous referee, Sébastien Lambert, Leonid Petrov, Jun Yang, Xiaopeng Cheng and Bo Zhang for the fruitful discussions and useful suggestions, which have greatly improved the work. This work was supported by the National Natural Science Foundation of China (NSFC) under grant Nos. 12573070 and 12373074. We acknowledge the use of the Radio Fundamental Catalogue (Petrov and Kovalev 2025). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations are managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research used material from the Bordeaux VLBI Image Database (BVID). This database can be reached at http://bvid.astrophy.u-bordeaux.fr/. We made much use of Astropy (http://www.astropy.org) – a community-developed core Python package for astronomy (2018AJ....156..123A), the Python 2D plotting library Matplotlib (2007CSE.....9...90H) and TOPCAT (2011ascl.soft01010T). This work benefited from the LaTeX service provided by the e-Science Center of the Collaborative Innovation Center of Advanced Microstructures, Nanjing University (https://sci.nju.edu.cn). Researchers interested in the sample-related data may contact us via email.
@@ Appendix B Cross-checking of redshifts from different methods
As shown in Fig. 9, we performed cross-checking of redshifts between different methods. DESI spectroscopic redshifts are compared with the literature-based redshifts (panel a), and SDSS spectroscopic redshifts from DR17 and DR19 (panel c). Similarly, SDSS spectroscopic redshifts are cross-checked against the literature-based redshifts (panel b). We find that the majority of sources show good agreement in redshifts across different methods ($|\Delta z|\!<\!0.1$). However, we find that DESI spectroscopic redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 show significant deviations from redshifts measured by other methods (panels a and c). In this work, for sources with both SDSS and DESI spectroscopic redshifts, we retained only those with consistent redshifts satisfying $|\Delta z|\!<\!0.1$ (panel d).
@@ Appendix C Astrometric properties as a function of $G$ magnitude
In this work, we find that the absolute radio-optical offsets (Fig. 10), the Gaia optical positional uncertainties (Fig. 11), parallax uncertainties (Fig. 12), proper motions and proper motion uncertainties (Fig. 13) in our sample all increase steeply toward fainter $G$ magnitudes and can be well described by an exponential dependence on $G$. In contrast, as shown in Fig. 11, the median radio positional uncertainties increase only mildly and approximately linearly with $G$ magnitude, from about 0.5 mas at $G\!\sim\!18$ to roughly 1.5 mas at $G\!\sim\!21$, which is significantly smaller than the corresponding optical positional uncertainties at faint $G$ magnitudes. However, at a given $G$ magnitude, both the absolute and relative radio–optical offsets are significantly larger for galaxies than for QSOs. The Gaia optical positional uncertainties are also significantly larger for galaxies than for QSOs, while the radio positional uncertainties are comparable between galaxies and QSOs. Taken together, these results indicate that the observed radio–optical offsets of RFC sources are most likely dominated by uncertainties in the Gaia optical positions rather than in the VLBI radio measurements.